Hi, I could use some help with this literal equation.
Solve for T: n = aT^2-4T+m
Thank you

Answer by Theo(13342)

start with:

n = aT^2 - 4T + m

subtract m from both sides of the equation to get:

n - m = aT^2 - 4T

factor out the a on the right side of the equation to get:

n - m = a * (T^2 - 4/a * T)

in the preceding step you had to be a little creative because a * -4/a*T = -4*T

divide both sides of the equation by a to get:

(n-m)/a = T^2 - 4/a * T)

complete the squares on the right side of the equation to get:

(n-m)/a = (T-2/a)^2 - 4/a^2

add 4/a^2 to both sides of the equation to get:

(n-m)/a + 4/a^2 = (T-2/a)^2

take the square root of both sides of the equation to get:

sqrt((n-m)/a + 4/a^2) = T - 2/a

add 2/a to both sides of the equation to get:

sqrt((n-m)/a - 4/a^2) + 2/a = T

that's your solution.

T = sqrt((n-m)/a + 4/a^2) + 2/a

the way to check is to give a random value to a and T and m and solve for n in the original equation.

then use that value of n and the values of a and m previously chosen to solve for T in the final equation.

I did and the solution is confirmed as good.

I used:

a = 9
m = 5
t = 2

i solved for n using the original equation to get n = 33.

i then used:

a = 9
m = 5
n = 33

i solved for t using the final equation to get t = 2.

that confirmed the solution is good.